OpenStudy (anonymous):

Given the function f(x) = 2x^3 - 5x^2. Use the expression lim h-->0 f(x) f(x+h)- f(x)/h Please help have exams in the morning :(

OpenStudy (anonymous):

So your expression is what's called the "difference quotient." It's a geometric way of getting at the tangent to a polynomial like f(x) here.

OpenStudy (anonymous):

\[ \frac{f(h+x)-f(x)}{h}=\frac{2 (h+x)^3-5 (h+x)^2-2 x^3+5 x^2}{h}=\\\frac{2 h^3+6 h^2 x-5 h^2+6 h x^2-10 h x}{h}=2 h^2+6 h x-5 h+6 x^2-10 x \]

OpenStudy (anonymous):

What he said

OpenStudy (anonymous):

lol^

OpenStudy (anonymous):

the limit when h goes to zero is \[ -10 x +6 x^2 \]

OpenStudy (anonymous):

definition of derivative =P

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ontech does this make sense?

OpenStudy (anonymous):

what she did is for f(x+h) , she replaced all the x's with (x+h) in the function \[f(x)=2x^3-5x^2\]

OpenStudy (anonymous):

Pretty sure he is the guy on the right. :)

OpenStudy (anonymous):

left

OpenStudy (anonymous):

And he is very quick with the equation editor.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Thanks .. im really trying to figure it out here..

OpenStudy (anonymous):

Stil there @ontech I have a comment if you want...

OpenStudy (anonymous):

im still here.. sorry for being so slow.. Im trying to figure it out as much as I can.. I need to understnad this..

OpenStudy (anonymous):

Seems like you are. I would forget the cofactors and expand the basic functions They expand with cofactors from Pascal's Triangle (x+h)^2 = x^2 + 2xh + h^2 (x+h)^3 = x^3 + 3x^2h + 3xh^2 + h^3 (x+h)^4 = .. the coefficients will be 1 4 6 4 1 Don't add the cofactors (like the 2 in 2x^3 in your problem) until you have it all multiplied out.

OpenStudy (anonymous):

And of course the basic point here is that you will subtract the first term (that's f(x)) and divide by h, and then all the other terms disappear as h -> 0, and you have the derivative by the power rule. Which is the whole point.

OpenStudy (anonymous):

ahhh